Topological properties of a model for

reconstructing transparent 3D

images

Oberwolfach, january 2007

Giovanni Bellettini Università di Roma - Tor Vergata

(joint work with V. Beorchia and M. Paolini)

Topological properties of a model for reconstructing transparent 3D images – p.1/2

Motivation: recover 3D surfaces

[show animation ./anim1 (ctrl-F3)]

Topological properties of a model for reconstructing transparent 3D images – p.2/2